How can you predict if an ionic compound will have a further hydrolysis reaction? (what needs to be true of at least one of its ions?)
Answer #1

1. The salt of weak and strong acid/base have some residual property of strong acid/base.

2. You need to calculate its hydrolysis constant, if it has some appreciable value then hydrolysis will occur else not

For example

CH3COOH + NaOH -----> CH3COONa + H2O

here CH3COONa is salt of strong base and weak acid and it has residual property of base and hence it will hydrolyse

Ka of CH3COOH = 1.8 x 18-5

hydrolysis constant Kh = Kw/Ka = 10-14/1.8 x 10-5 = 5.55 x 10-10

If you see salt of HCl and NaOH both are strong and its salt won't hydrolyse

If you calculate its Kh it will be very - very less than 1 , of the order of 10-30 and hence not much appreciable hydrolysis